Fermi National Laboratory
 
Operation Summary for March 19, 2012 - March 21, 2012
 

Monday March 19

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI, MiniBooNE, SeaQuest, and FTBFT-1017 experiments taking beam.

During the day shift, Muon Ring personnel began a study using the antiproton target to create muons at 9:01 AM. At 9:41 AM, a Booster bulk power supply (B:CEP400) tripped off; it reset. The NuMI beam permit tripped off at 10:56 AM on an incomplete RF gap indication, but no MI RF station had tripped off; the beam permit reset. At 11:11 AM, the NuMI beam permit tripped off due to the output of a power supply (E:V118) drifting low; everything reset. The Booster West Anode Power Supply tripped off at 11:23 AM; it reset. The Muon Ring study ended at 12:08 PM. A MI person began a g-2 (g minus 2) study at 1:09 PM. The MI RF South Anode Power Supply tripped off on an interlock fault indication; everything reset. T-1017 and SeaQuest made accesses during this shift.

During the evening shift, the MI g-2 study ended at 6:15 PM. The NuMI beam permit tripped off at 10:31 PM due to V118; it reset.

Tuesday March 20

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1017.

During the midnight shift, Operations completed a 750 KeV line tune at 3:04 AM. Operations reported tuning the Booster notch at 4:27 AM. At 4:46 AM, a Linac RF station (LRF1) tripped off; Operators had in back up by 5:06 AM. At 6:54 AM, the NuMI beam permit tripped off due to V118; everything reset. Operations completed a 400 MeV line steer and a Boosted injection tune at 7:36 AM.

During the day shift, a Muon Ring person began a study around 8:59 AM. At 9:43 AM, a MiniBooNE person implemented a new run mode. At 12:30 PM, the NuMI beam permit tripped off due to V118.

During the evening shift, the Muon Ring study ended at 4:39 PM. At 6:35 PM, LRF1 tripped off and would not reset; Operators called in Linac personnel and entered an energy conservation timeline into the Time Line Generator (TLG). Linac personnel had LRF1 back online and Linac resumed sending beam at 7:09 PM. At 8:17 PM, the NuMI beam permit tripped off due to V118 for the first of three times in about three hours; it reset. During this shift, SeaQuest took a 20-minute run of high intensity pulses – Operations turned off T-1017 for this run.

Wednesday March 21

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1017.

During the midnight shift, Operations adjusted the 400 MeV line Debuncher cavity phase set point at 5:57 AM, which helped the Booster efficiency. The NuMI beam permit tripped off at 7:03 AM due to V118; it reset. Operations completed a Booster injection tune at 7:26 AM.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo agrees with the Booster and MI study plans for today. The MI study will halt beam to T-1017. SeaQuest began a full shift access for alignment and repairs.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/21/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory